OpenStudy (anonymous):

Solve the equation for w w - 5/3 = w/6

OpenStudy (anonymous):

can someone please help me with this one

OpenStudy (anonymous):

Try to multiply the whole equation by 6. That means both terms on the left side. Now you have 6W-30/3=W Now rearrange so all the w's are on one side

OpenStudy (anonymous):

w-5/3=w/6 +5/3 -w/6 both sides ------------ w-w/6=5/3 w(1-(1/6))=5/3 divide both sides by () w=(5/3)/(1-(1/6))

OpenStudy (anonymous):

w(1-1/6) = 5/3 =>5w=5x6/3 => w = 2